A battery of emf 100 V and a resistor of resistance 10 kΩ are joined in series. This system is used as a source to supply current ot an external resistance R. If R is not greater than 100 Ω, the current through it is constant upto two significant digits. Find its value. This is the basic principle of a constant-current source.

Grade:11

1 Answers

Kevin Nash
9 years ago
Sol. E = 100 V, R’ = 100 kΩ = 100000 Ω R = 1 – 100 When no other resister is added or R = 0. i = E/R’ = 100/100000 = 0.00 1Amp when R = 1 i = 100/100000 + 1 = 100/100001 = 0.0009A when R = 100 i = 100/100000 + 1 = 100/100100 = 0.0009999 A. Upto R = 100 the current does not upto 2 significant digits. Thus it proved.